How much water is needed to make a 1M solution of NaCl that contains 0.5moles
1
Expert's answer
2016-05-11T08:42:03-0400
CM(NaCl)=n(NaCl)/V(NaCl)
V=n/ CM
As NaCl is used in a dry form, the volume of solution will be equal to the volume of water.
V(H2O)=0.5/ 1=1 l